OpenStudy (anonymous):

Solve for x, given the equation √(x-5)+7=11

OpenStudy (the_fizicx99):

Subtract 7 from both sides, √(x-5) + 7 = 11 -7 -7 √(x-5) = 4 Square both sides, reason why you're squaring both sides it because you want to get rid of the radical, x-5 = 16 +5 +5 x = 21 Check, if it doesn't work its an extraneous solution meaning it doesn't work with the original equation. √(21-5) + 7 = 11 √(16) + 7 = 11 4 + 7 = 11 11 = 11 therefore this solution works

OpenStudy (anonymous):

Trapezoid JKLM is shown on the coordinate plane below. Trapezoid JKLM on the coordinate plane with ordered pairs at J negative 7, negative 2, at K negative 4, negative 2, at L negative 2, negative 5, at M negative 9, negative 5. If trapezoid JKLM is translated according to the rule (x, y) → (x + 8, y - 3), what are the coordinates of point L'? (1, -5) (-10, 6) (-5, 3) (6, -8)